an internal study at lahey electronics ,a large software developer company ,revealed the mean time for an internal e-mail message to arrive at its destination was 2 seconds. Further , the distribution of the arrival times followed the poisson distribution .what is the probability a message takes exactly 1 second to arrive at its destination?
Answer by Boreal(15235)

P(1)=e^(-2)*2^1/1!
Population parameter is 2
P(x)=e^(-x)x^n/n!, where x=2
that probability is 2 e^(-2)=0.2707